Two small spheres spread 10 cm apart have equal but opposite charge. If the magnitude of the force between them is 2.5 x 10^-10 N
(a)Calculate the charge on each sphere.
(b)Calculate the electric field intensity at the mid-point.
(c)Calculate the electric potential at the mid-point.
(d) How many excess electrons must be present on each sphere.
1
Expert's answer
2019-06-25T09:16:54-0400

a) According to Coulomb's law for two equal charges:


"F=\frac{1}{4\pi\epsilon_0}\cdot\frac{q^2}{r^2},"

"q=2r\sqrt{\pi\epsilon_0F}=1.667\cdot10^{-11}\text{ C}."

b) The field at the midpoint is a sum of fields of two charges. Since they have opposite signs, in the middle the field is twice as the field of a single charge (because the field is a vector quantity):


"E=2\cdot\frac{1}{4\pi\epsilon_0}\cdot\frac{q}{(r/2)^2}=119.86\text{ V/m}."

c) The electric potential is 0 in the middle since the two charges are equal but opposite:


"V=\frac{1}{4\pi\epsilon_0}\cdot\frac{q}{r/2}-\frac{1}{4\pi\epsilon_0}\cdot\frac{q}{r/2}=0."

d) It's easy to calculate by dividing the total charge of a sphere by the charge of an electron:


"N=\frac{q}{e}=104050000."